# Semiclassical descriptions of dissipative dynamics of strongly interacting Bose gases in optical lattices

### v2

@@ Appendix C Higher-order corrections and differential equations in the dTWA
Let us introduce one-site reduced operators ${\mathscr{A}}_{j}^{(1)}(t)={\rm Tr}_{\underline{j}}{\mathscr{A}}^{\otimes N}_{t}$ for the phase point operator acting on the whole system. Here, $\underline{j}$ denotes a trace operation over all sites except for site $j$. In parallel to this, two-site reduced operators, three-site reduced operators, and higher rank reduced operators are successively defined as ${\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)={\rm Tr}_{{\underline{j}},{\underline{j}}^{\prime}}{\mathscr{A}}^{\otimes N}_{t}$, ${\mathscr{A}}_{j,j^{\prime},j^{\prime\prime}}^{(3)}(t)={\rm Tr}_{{\underline{j}},{\underline{j}}^{\prime},{\underline{j}}^{\prime\prime}}{\mathscr{A}}^{\otimes N}_{t}$, and so on. The time evolution equation of ${\mathscr{A}}_{j}^{(1)}(t)$ is derived by tracing the $N-1$ site indices in Eq. (9). However, it is not a closed equation for ${\mathscr{A}}_{j}^{(1)}(t)$ itself. Indeed, some inseparable correlations in ${\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)$, i.e., ${\mathscr{B}}_{j,j^{\prime}}^{(2)}(t)\equiv{\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)-{\mathscr{A}}_{j}^{(1)}(t){\mathscr{A}}_{j^{\prime}}^{(1)}(t)$, may appear in the determination. Similarly, to update the values of ${\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)$, the information of ${\mathscr{A}}_{j,j^{\prime},j^{\prime\prime}}^{(3)}(t)$ is necessary, and more generally, updating ${\mathscr{A}}_{j,j^{\prime},\cdots}^{(s-1)}(t)$ requires ${\mathscr{A}}_{j,j^{\prime},j^{\prime\prime},\cdots}^{(s)}(t)$ at the past step. This hierarchy is known as the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy in the dynamics of multipoint correlation functions Pucci et al. (2016).
@@
The dTWA is formulated as a first order truncation method of this hierarchy within the single-site decoupling approximation. In this approximation, each two-site operator is approximated to a factorized operator ${\mathscr{A}}^{(2)}_{j,j^{\prime}}(t)={\mathscr{A}}^{(1)}_{j}(t){\mathscr{A}}^{(1)}_{j^{\prime}}(t)$, i.e., ${\mathscr{B}}_{j,j^{\prime}}^{(2)}(t)=0$. Any higher-order operator also factorizes. Hence, the time evolution of the one-site operators can be determined in a closed manner. This hierarchy truncation is equivalent to the ansatz assumed in Sec. III.4. A single-site update of the state is performed by integrating the classical equation of motion for single-site variables for a finite interval of time. Beyond this leading-order treatment, it is also possible to incorporate inseparable multisite reduced operators step by step in an explicit manner. However, the computational cost to handle the corresponding time evolution equations increases exponentially as the number of indices involved in an inseparable operator increases.
@@
Notice that, reflecting the trace-conserving property of the time evolution generator in the master equation, which implies that $\langle{\hat{1}}\rangle(t)=1$ holds for all time, the coefficient in front of $I$ remains constant in ${\mathscr{A}}_{j,{\rm dTWA}}^{(1)}(t)$. However, this is not necessarily the case if the non-unitary time evolution belongs to a specific class, such as non-unitary dynamics for a non-Hermitian Hamiltonian, which violates the normalization of the wave functions. A substantial discussion on this issue can be found in Ref. Singh and Weimer (2022), where the authors attempt to solve the dTWA equation for the single-site operator by adopting the quantum jump method, which requires semiclassical evaluations of quantum trajectories produced by a non-Hermitian Hamiltonian.
@@
Here, the derivatives of the Hamiltonian read
@@ Appendix D Numerical sampling scheme for phase-space trajectories
As described in Sec. III.4, to adequately represent quantum fluctuations embedded in the unit filling Mott insulating state $|\psi_{\rm Mott}\rangle=\prod_{j}{\hat{a}}^{\dagger}_{j}|0\rangle=\prod_{j}|1\rangle_{j}$, we use the GDW sampling scheme for SU(3) matrices Zhu et al. (2019); Nagao et al. (2021). This sampling scheme is based on a projective decomposition of each SU(3) base matrix, i.e.,
@@
where $\lambda^{(a)}_{\alpha}$ is the $a$th eigenvalue of $T_{\alpha}$ and $|\phi_{\alpha}^{(a)}\rangle$ is the corresponding eigenstate. The probability to find ${\hat{X}}^{(j)}_{\alpha}$ in $\lambda^{(a)}_{\alpha}$ is given by $p_{\alpha,a}^{(j)}=|\langle\phi^{(a)}_{\alpha}|1\rangle_{j}|^{2}\geq 0$. Since $T_{3}$ and $T_{8}$ are already diagonal in our definition (see Appendix B), the classical variables $r_{3}^{(j)}$ and $r_{8}^{(j)}$ take 00 and $2/\sqrt{3}$, respectively, with a probability of $1$. In addition, $r_{4}^{(j)}$ and $r_{5}^{(j)}$, which correspond to nondiagonal matrices, are zero also with a probability of $1$. However, the other variables, i.e., $r_{1}^{(j)}$, $r_{2}^{(j)}$, $r_{6}^{(j)}$, and $r_{7}^{(j)}$, discretely fluctuate over binary values in $\{-1,+1\}$ with an equal probability of $1/2$. Therefore, the local joint probability distribution for the variables, which provides the proper discrete Wigner function for the local Fock vector, reads ${\mathscr{W}}^{[j]}_{1}(r_{1}^{(j)},\cdots,r_{8}^{(j)})=\delta(r_{3}^{(j)})\delta(r_{4}^{(j)})\delta(r_{5}^{(j)})\delta(r_{8}^{(j)}-2/\sqrt{3})\prod_{s=1,2,6,7}[\frac{1}{2}\delta(r_{s}^{(j)}-1)+\frac{1}{2}\delta(r_{s}^{(j)}+1)]$. We can easily check that this probability distribution gives the correct coefficient in the expansion of $|1\rangle\langle 1|$ with respect to ${\mathscr{A}}^{[j]}(r_{1}^{(j)},\cdots,r_{8}^{(j)})$. A physical understanding of this probability distribution is that the fluctuations of the four variables describe the phase fluctuations in the Mott insulating phase.
@@ Appendix E Details of the Fokker-Planck TWA for the SU(3) variables
Here, the differential operator, which can be referred to as a Bopp operator, acts on the Wigner function $W_{\rm eff}\equiv({\hat{\rho}}_{\rm eff})_{W}$, defined as the continuous Weyl symbol of ${\hat{\rho}}_{\rm eff}$. The first term in Eq. (28) gives the classical limit of the operator ${\hat{X}}^{(j)}_{\alpha}$, i.e., the Weyl symbol of the SU(3) spin operator given as $x_{\alpha}^{(j)}=({\hat{X}}_{\alpha}^{(j)})_{W}$. The second term in Eq. (28), in the form of a differential operator, emerges beyond the classical limit to restore the noncommutativity between the spin operators. Indeed, Eq. (28) implies that the commutator between ${\hat{X}}_{\alpha}^{(j)}$ and ${\hat{X}}_{\beta}^{(j)}$ is transformed as $[{\hat{X}}_{\alpha}^{(j)},{\hat{X}}_{\beta}^{(j)}]\rightarrow if_{\alpha\beta\gamma}x_{\gamma}^{(j)}=\{x_{\alpha}^{(j)},x_{\beta}^{(j)}\}_{\rm P.B.}$, and hence the amount of noncommutativity is quantified in the phase space by the corresponding Poisson bracket for the classical variables. The Bopp operator given above ignores higher-order derivatives beyond the leading order contribution of the first-order derivatives, thus allowing one to readily derive the SU(3) TWA for open systems as well as isolated systems.
@@
Notice that $f_{{\rm PA},\alpha}^{(j)}$ are nonlinear functions of the phase space variables, implying the limitation of the leading order truncation in the Bopp operators.
We now explain how the numerical sampling is carried out to compute an average over stochastic Langevin trajectories. First, we draw a set of random spins from a Wigner distribution. In this study, we use a generalized discrete Wigner sampling scheme, described in Appendix D, for the purpose of sampling. The initialized spins are then propagated with the Langevin equation in Eq. (14) until the final time at $t=t_{1}>0$, so that we have a non-differentiable trajectory $x^{(j)}_{{\rm cl},\alpha}(t_{1})=x^{(j)}_{{\rm cl},\alpha}[t_{1};x^{(j)}_{{\rm cl},\alpha}(n\Delta t),\cdots,x^{(j)}_{{\rm cl},\alpha}(\Delta t),x^{(j)}_{{\rm cl},\alpha}(0)]$. To numerically integrate the Langevin equation, we employ the first-order Euler-Maruyama algorithm Gardiner (2009) with $\Delta t=t_{1}/(n+1)$, where $n$ is an integer providing a sufficiently small $\Delta t$, ensuring the stability of the simulation. This trajectory, also referred to as a path of the Wiener process, is a function of generated random numbers $dw_{q}^{(j)}(t)$ during the time sequence and is thus not differentiable everywhere. Using the registered data of this trajectory, we evaluate ${\cal O}_{W}\left[\{x^{(j)}_{{\rm cl},\alpha}(t)\}\right]$ for each $t\in[0,t_{1}]$, which is a fluctuating classical function. We iteratively generate many stochastic trajectories and take the Monte Carlo average of ${\cal O}_{W}\left[\{x^{(j)}_{{\rm cl},\alpha}(t)\}\right]$ over these trajectories until the convergence is finally reached. In contrast to the dTWA case, the FP-TWA must successively generate Gaussian noises at each time step after taking initial noises at $t=0$ from the Wigner distribution. We should also note that higher order algorithms for solving the Langevin equation, such as the Platen scheme to second order of $\Delta t$ Breuer and Petruccione (2002), are quite hard to implement in practice due to the non-differentiable manner of trajectories.

### v3

@@ Appendix C Higher-order corrections and differential equations in the dTWA
Let us introduce one-site reduced operators ${\mathscr{A}}_{j}^{(1)}(t)={\rm Tr}_{\underline{j}}{\mathscr{A}}^{\otimes N}_{t}$ for the phase-point operator acting on the whole system. Here, $\underline{j}$ denotes a trace operation over all sites except for site $j$. In parallel to this, two-site reduced operators, three-site reduced operators, and higher-rank reduced operators are successively defined as ${\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)={\rm Tr}_{{\underline{j}},{\underline{j}}^{\prime}}{\mathscr{A}}^{\otimes N}_{t}$, ${\mathscr{A}}_{j,j^{\prime},j^{\prime\prime}}^{(3)}(t)={\rm Tr}_{{\underline{j}},{\underline{j}}^{\prime},{\underline{j}}^{\prime\prime}}{\mathscr{A}}^{\otimes N}_{t}$, and so on. The time-evolution equation of ${\mathscr{A}}_{j}^{(1)}(t)$ is derived by tracing the $N-1$ site indices in Eq. (9). However, it is not a closed equation for ${\mathscr{A}}_{j}^{(1)}(t)$ itself. Indeed, some inseparable correlations in ${\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)$, i.e., ${\mathscr{B}}_{j,j^{\prime}}^{(2)}(t)\equiv{\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)-{\mathscr{A}}_{j}^{(1)}(t){\mathscr{A}}_{j^{\prime}}^{(1)}(t)$, may appear in the determination. Similarly, to update the values of ${\mathscr{A}}_{j,j^{\prime}}^{(2)}(t)$, the information of ${\mathscr{A}}_{j,j^{\prime},j^{\prime\prime}}^{(3)}(t)$ is necessary, and more generally, updating ${\mathscr{A}}_{j,j^{\prime},\cdots}^{(s-1)}(t)$ requires ${\mathscr{A}}_{j,j^{\prime},j^{\prime\prime},\cdots}^{(s)}(t)$ at the past step. This hierarchy is known as the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy in the dynamics of multipoint correlation functions Pucci et al. (2016).
@@
The dTWA is formulated as a first-order truncation method of this hierarchy within the single-site decoupling approximation. In this approximation, each two-site operator is approximated to a factorized operator ${\mathscr{A}}^{(2)}_{j,j^{\prime}}(t)={\mathscr{A}}^{(1)}_{j}(t){\mathscr{A}}^{(1)}_{j^{\prime}}(t)$, i.e., ${\mathscr{B}}_{j,j^{\prime}}^{(2)}(t)=0$. Any higher-order operator also factorizes. Hence, the time evolution of the one-site operators can be determined in a closed manner. This hierarchy truncation is equivalent to the ansatz assumed in Sec. III.4. A single-site update of the state is performed by integrating the classical equation of motion for single-site variables for a finite interval of time. Beyond this leading-order treatment, it is also possible to incorporate inseparable multisite reduced operators step by step in an explicit manner. However, the computational cost to handle the corresponding time-evolution equations increases exponentially as the number of indices involved in an inseparable operator increases.
@@
Notice that, reflecting the trace-conserving property of the time-evolution generator in the master equation, which implies that $\langle{\hat{1}}\rangle(t)=1$ holds for all time, the coefficient in front of $I$ remains constant in ${\mathscr{A}}_{j,{\rm dTWA}}^{(1)}(t)$. However, this is not necessarily the case if the nonunitary time evolution belongs to a specific class, such as nonunitary dynamics for a non-Hermitian Hamiltonian, which violates the normalization of the wave functions. A substantial discussion on this issue can be found in Ref. Singh and Weimer (2022), where the authors attempt to solve the dTWA equation for the single-site operator by adopting the quantum jump method, which requires semiclassical evaluations of quantum trajectories produced by a non-Hermitian Hamiltonian.
@@
Here, the derivatives of the Hamiltonian read as
@@ Appendix D Numerical sampling scheme for phase-space trajectories
As described in Sec. III.4, to adequately represent quantum fluctuations embedded in the unit filling Mott-insulating state $|\psi_{\rm Mott}\rangle=\prod_{j}{\hat{a}}^{\dagger}_{j}|0\rangle=\prod_{j}|1\rangle_{j}$, we use the GDW sampling scheme for SU(3) matrices Zhu et al. (2019); Nagao et al. (2021). This sampling scheme is based on a projective decomposition of each SU(3) base matrix, i.e.,
@@
where $\lambda^{(a)}_{\alpha}$ is the $a$th eigenvalue of $T_{\alpha}$ and $|\phi_{\alpha}^{(a)}\rangle$ is the corresponding eigenstate. The probability to find ${\hat{X}}^{(j)}_{\alpha}$ in $\lambda^{(a)}_{\alpha}$ is given by $p_{\alpha,a}^{(j)}=|\langle\phi^{(a)}_{\alpha}|1\rangle_{j}|^{2}\geq 0$. Since $T_{3}$ and $T_{8}$ are already diagonal in our definition (see Appendix B), the classical variables $r_{3}^{(j)}$ and $r_{8}^{(j)}$ take 00 and $2/\sqrt{3}$, respectively, with a probability of $1$. In addition, $r_{4}^{(j)}$ and $r_{5}^{(j)}$, which correspond to nondiagonal matrices, are zero also with a probability of $1$. However, the other variables, i.e., $r_{1}^{(j)}$, $r_{2}^{(j)}$, $r_{6}^{(j)}$, and $r_{7}^{(j)}$, discretely fluctuate over binary values in $\{-1,+1\}$ with an equal probability of $1/2$. Therefore, the local joint probability distribution for the variables, which provides the proper discrete Wigner function for the local Fock vector, reads ${\mathscr{W}}^{[j]}_{1}(r_{1}^{(j)},\cdots,r_{8}^{(j)})=\delta(r_{3}^{(j)})\delta(r_{4}^{(j)})\delta(r_{5}^{(j)})\delta(r_{8}^{(j)}-2/\sqrt{3})\prod_{s=1,2,6,7}[\frac{1}{2}\delta(r_{s}^{(j)}-1)+\frac{1}{2}\delta(r_{s}^{(j)}+1)]$. We can easily check that this probability distribution gives the correct coefficient in the expansion of $|1\rangle\langle 1|$ with respect to ${\mathscr{A}}^{[j]}(r_{1}^{(j)},\cdots,r_{8}^{(j)})$. A physical understanding of this probability distribution is that the fluctuations of the four variables describe the phase fluctuations in the Mott-insulating phase.
@@ Appendix E Details of the Fokker-Planck TWA for the SU(3) variables
Here, the differential operator, which can be referred to as a Bopp operator, acts on the Wigner function $W_{\rm eff}\equiv({\hat{\rho}}_{\rm eff})_{W}$, defined as the continuous Weyl symbol of ${\hat{\rho}}_{\rm eff}$. The first term in Eq. (28) gives the classical limit of the operator ${\hat{X}}^{(j)}_{\alpha}$, i.e., the Weyl symbol of the SU(3) spin operator given as $x_{\alpha}^{(j)}=({\hat{X}}_{\alpha}^{(j)})_{W}$. The second term in Eq. (28), in the form of a differential operator, emerges beyond the classical limit to restore the noncommutativity between the spin operators. Indeed, Eq. (28) implies that the commutator between ${\hat{X}}_{\alpha}^{(j)}$ and ${\hat{X}}_{\beta}^{(j)}$ is transformed as $[{\hat{X}}_{\alpha}^{(j)},{\hat{X}}_{\beta}^{(j)}]\rightarrow if_{\alpha\beta\gamma}x_{\gamma}^{(j)}=\{x_{\alpha}^{(j)},x_{\beta}^{(j)}\}_{\rm P.B.}$, and hence the amount of noncommutativity is quantified in the phase space by the corresponding Poisson bracket for the classical variables. The Bopp operator given above ignores higher-order derivatives beyond the leading-order contribution of the first-order derivatives, thus allowing one to readily derive the SU(3) TWA for open systems as well as isolated systems.
@@
Notice that $f_{{\rm PA},\alpha}^{(j)}$ are nonlinear functions of the phase-space variables, implying the limitation of the leading order truncation in the Bopp operators.
We now explain how the numerical sampling is carried out to compute an average over stochastic Langevin trajectories. First, we draw a set of random spins from a Wigner distribution. In this study, we use a generalized discrete Wigner sampling scheme, described in Appendix D, for the purpose of sampling. The initialized spins are then propagated with the Langevin equation in Eq. (14) until the final time at $t=t_{1}>0$, so that we have a non-differentiable trajectory $x^{(j)}_{{\rm cl},\alpha}(t_{1})=x^{(j)}_{{\rm cl},\alpha}[t_{1};x^{(j)}_{{\rm cl},\alpha}(n\Delta t),\cdots,x^{(j)}_{{\rm cl},\alpha}(\Delta t),x^{(j)}_{{\rm cl},\alpha}(0)]$. To numerically integrate the Langevin equation, we employ the first-order Euler-Maruyama algorithm Gardiner (2009) with $\Delta t=t_{1}/(n+1)$, where $n$ is an integer providing a sufficiently small $\Delta t$, ensuring the stability of the simulation. This trajectory, also referred to as a path of the Wiener process, is a function of generated random numbers $dw_{q}^{(j)}(t)$ during the time sequence and is thus not differentiable everywhere. Using the registered data of this trajectory, we evaluate ${\cal O}_{W}\left[\{x^{(j)}_{{\rm cl},\alpha}(t)\}\right]$ for each $t\in[0,t_{1}]$, which is a fluctuating classical function. We iteratively generate many stochastic trajectories and take the Monte Carlo average of ${\cal O}_{W}\left[\{x^{(j)}_{{\rm cl},\alpha}(t)\}\right]$ over these trajectories until the convergence is finally reached. In contrast to the dTWA case, the FP-TWA must successively generate Gaussian noises at each time step after taking initial noises at $t=0$ from the Wigner distribution. We should also note that higher-order algorithms for solving the Langevin equation, such as the Platen scheme to second order of $\Delta t$ Breuer and Petruccione (2002), are quite hard to implement in practice due to the nondifferentiable manner of trajectories.